OpenStudy (anonymous):

Write the equation of the line that passes through (4, 2) and is parallel to the line y = 2x – 1. A. x – y = 6 B. 2x – y = 6 C. 2x – y = 1 D. 2x – y = –1

OpenStudy (anonymous):

y = 2x - 1 can you tell me the slope of this line ?

OpenStudy (anonymous):

do you know how to find the slope ?

OpenStudy (anonymous):

In y = mx + b form (which is what your equation is in), the slope is in the m position. So can you tell me the slope ?

OpenStudy (anonymous):

I will do this one for you, just to show you. y = mx + b y = 2x - 1 the number in the m position is 2, therefore the slope is 2. A parallel line will have the same slope. Now we will go back to y = mx + b slope(m) = 2 (4,2) .. x = 4 and y = 2 now we will sub our info into the formula and solve for b (the y intercept) 2 = 2(4) + b 2 = 8 + b 2 - 8 = b -6 = b equation of parallel line is : y = 2x - 6, but it appears you need it in standard form : Ax + By = C y = 2x - 6 -- subtract 2x from both sides -2x + y = -6 -- multiply by -1 to make x positive 2x - y = 6 <== your parallel line in standard form

OpenStudy (anonymous):

thanks sweetie <3 #Medal